# http://dreamscene360.github.io/

### v1

@@ 4.3 Ablation Study
5 Conclusion and Limitations
In this work, we present DreamScene360, a novel framework that unrestrictedly generates high-quality, immersive 3D scenes with complete ${}^{\circ}$ coverage in various styles from text inputs, regardless of their specificity level. Our pipeline leverages panorama as a middle ground, which provides us with a self-refinement opportunity by asking GPT-4V for quality assessment and prompt revision. Subsequently, we construct a geometric field that initializes 3D Gaussians. With the help of unsupervised geometric and semantic loss on virtual views, we lift the 2D panorama into panoramic 3D Gaussians. Experiments demonstrate our superiority against baseline methods in terms of global consistency and visual quality.
Limitations
Despite the exciting results presented by DreamScene360, our generation results are still limited at the resolution of the text-to-image diffusion model ($512\times 1024$). Moving forward, we will explore generating 3D scenes at higher resolution for a better, seamless, immersive user experience.

### v2

@@ 4.3 Ablation Study
5 Conclusion
In this work, we present DreamScene360, a novel framework that unrestrictedly generates high-quality, immersive 3D scenes with complete 360 coverage in various styles from text inputs, regardless of their specificity level. Our pipeline leverages panorama as a middle ground, which provides us with a self-refinement opportunity by asking GPT-4V for quality assessment and prompt revision. Subsequently, we construct a geometric field that initializes 3D Gaussians. With the help of unsupervised geometric and semantic loss on virtual views, we lift the 2D panorama into panoramic 3D Gaussians. Experiments demonstrate our superiority against baseline methods in terms of global consistency and visual quality. However, our generation results are still limited by the default resolution ($512\times 1024$) of the pretrained text-to-panorama diffusion model. Moving forward, we will explore generating 3D scenes at higher resolutions and even 4D dynamic scenes [28] for a better, seamless, and immersive user experience.
Acknowledgements
We thank the Visual Machines Group (VMG) at UCLA, Visual Informatics Group at UT Austin (VITA), and ARL Intelligent Perception Branch for feedback and support. This project was supported by LUCI program under the Basic Research Office and partially supported by ARL grants W911NF20-2-0158 and W911NF-21-2-0104 under the cooperative A2I2 program. Z.W. is supported by the U.S. DEVCOM Army Research Laboratory and LUCI program, and an Army Young Investigator Award. A.K. is supported by a DARPA Young Faculty Award, NSF CAREER Award IIS-2046737, and Army Young Investigator Award.